OpenStudy (liziekay):

A bullet is fired into the air with an initial velocity of 1,200 feet per second at an angle for 45° from the horizontal. What is the horizontal distance traveled by the bullet in 2 seconds? (Neglect the resistance of air on the bullet. Round your answer to the nearest whole number.) ____ ft. PLEASE HELP!

OpenStudy (liziekay):

@pooja195

OpenStudy (brainzonly):

you've been waiting a while...

OpenStudy (brainzonly):

I have no idea about this one, it might take me a while to get an answer.

OpenStudy (brainzonly):

@mathmale A bullet is fired into the air with an initial velocity of 1,200 feet per second at an angle for 45° from the horizontal. What is the horizontal distance traveled by the bullet in 2 seconds? (Neglect the resistance of air on the bullet. Round your answer to the nearest whole number.) ____ ft.

OpenStudy (holsteremission):

First find the horizontal component of the initial velocity vector, then multiply that velocity by the given time.

OpenStudy (liziekay):

@HolsterEmission Can you show me an example? Like a diagram maybe?

OpenStudy (holsteremission):

Sure: Basic trig tells you that \(\cos\theta=\dfrac{\|\vec{v}_x\|}{\|\vec{v}\|}\), or \(\|\vec{v}_x\|=\|\vec{v}\|\cos\theta\). (Here \(\|\cdot\|\) means magnitude.) You're given that \(\|\vec{v}\|=1200\) and \(\theta=45^\circ=\dfrac{\pi}{4}\text{ rad}\), so it's just a matter of plugging in these values.